# Active Sensing of Knee Osteoarthritis Progression with Reinforcement Learning

### v1

@@ 3.1 Knee osteoarthritis
In the KOA research area, there has been a growing interest in the application of machine learning (ML) methods for KOA progression prediction [74, 77, 23, 43, 32, 76, 51, 49, 54, 55, 13, 27]. Two main lines of work related to the prediction of KOA progression from clinical and imaging data are the prediction of TKR [77, 43, 32] and prediction of radiographic KOA worsening [74, 23, 49].
@@
The main limitation of most previous studies on KOA progression prediction is that they consider predictions per knee, disregarding that OA often co-occurs in two knees. The second limitation is that they ignore the long disease progression timelines, taking for modeling only one baseline measurement point. OA can progress very slowly, and be drastically accelerated by various factors, e.g. injury occurring at some point in the future. Therefore, we believe that it is crucial to update the model’s predictions when newer data from a patient becomes available. Finally, none of the previous studies considered the financial aspect of their results, nor downstream drug development or clinical applications. Our work aims to address all these limitations, enabling dynamic patient-level data collection, while balancing data collection utility and costs with RL.
@@ 3.2 Active sensing
RL has become one of the most common techniques to explore AS in recent years, particularly in medical applications for disease management [2, 83, 10, 62, 29]. DPSCREEN framework proposed by [2] aimed to deliver a personalized screening policy for breast cancer patients. The proposed policy takes into account both patient features and their clinical history, and it demonstrated significant reductions in the number of screenings performed, while maintaining the same expected delays in disease detection. [83] performed the AS using the Actor-Critic (ASAC) framework to tackle the issue of selecting the variables that should be observed. The ASAC framework includes two networks: a selector network, playing as an actor, and a predictor network, playing as a critic. The authors showed that ASAC outperformed state-of-the-art methods in two real-world medical datasets of Alzheimer’s Disease and electronic health records. [10] presented a deep Q-learning approach to strategically schedule measurements, especially in ICU mortality prediction. In simulations and real-world datasets, the proposed policy outperforms heuristic-based scheduling with higher predictive gain and lower cost.
The prior work on AS with RL focused on improving general RL methods and placed a substantially lesser focus on the reward function, the disease of interest, as well as the challenge of handling high-dimensional data. In this work, we chose a simple RL technique – Q-learning, and instead focused on the development of the reward and the environment. We believe that our work lays the foundations for future, more RL-oriented studies for AS.
@@ 4.2 Hyper-parameters of the reward function
We empirically chose $\alpha=0.5$ as a parameter to scale down the late visit-associated negative reward, and set $\beta=0.3c$ to assign positive rewards for true dismissal. $\lambda=0.5$ was used in all our experiments, and we also conducted an ablation study to validate how our method performs in different settings. More details about these hyper-parameters will be presented in Section 5.2.
@@ 4.3 Agent
[…]
4.4 Evaluation
We compared our method to a wide range of reference methods belonging to three categories of sensing – random, routine, and active sensing. Random policies combined $4$ probabilities of taking follow-up action: $0.3$, $0.5$, $0.7$, $0.9$. The set of routine policies comprised “no sensing” (NS), annual sensing (ANS), and bi-annual sensing (BAS). The active policies set included logistic regression (LR), cox regression (CR), RNN-based methods, and CLIMATv2 [51, 49]. The reference methods are described in detail in Section S1.1.
@@ 5.2 Impact of financial factors on decision-making
We demonstrated the second experiment in Figure 4b. The results from CLIMATv2 and the NS policy exhibited a decreasing trend as the JSW cost, $c$, increased while keeping unchanged $\lambda=0.5$. In contrast, we observed that the ANS policy generally increased in proportion to $c$. Our policy achieved a constant reward of $-0.25$ for $c\leq 1.0$, and demonstrated a linear increase for $c>1.0$. Overall, our proposed policy gained the highest reward among the methods at any value of $c\in[0.5,2]$. For values of $c$ greater than $1.5$, our method stands out as the sole approach yielding positive rewards per person.
@@ 5.3 Patient-level versus knee-level predictions
Table 4 demonstrates the overall performance of CLIMATv2 and our proposed method with knee-level and patient-level approaches. Notably, the patient-level approach outperformed the knee-level approach for both CLIMATv2 and the RL-based policy, with average improvements in BA of $1.44\%$ and $3.73\%$, respectively, and in RC of $4.58\%$ and $20.25\%$, respectively. Both the methods gained a $0.02$ increase in RPP when switching from the knee-level to the patient-level approach. To provide additional insights, we further present the corresponding fine-grained results in Figure 5. Most of the time, patient-level models reduced the standard error in BA and RC of both CLIMATv2 and our RL-based agent. Particularly, at the 1-year time point, our patient-level policy exceeded the patient-level CLIMATv2 with the highest difference by approximately 15% in BA and 60% in RC. Furthermore, the greatest gap between patient-level and knee-level of our method occurred 3 years after baseline by approximately 20% in BA and 40% in RC. In general, our patient-level policy showed substantial improvement compared to other methods most of the time.
@@ 5.4 Importance of imaging data
We also evaluated the impact of imaging data on the performance of the proposed method. To set up a benchmark, we conducted experiments without the contributions of radiographs, thereby excluding the KLG probabilities from the agent’s input. We then incorporated one of the three distinct types of KLGs into our experiments. The first type, referred to as the “KLG ground truth” corresponds to the manual grades produced by the radiologists. The second and third type named the “image features” and the “KLG probabilities”, respectively, were extracted from a DNN that had been trained to grade the KLGs automatically.
@@ 5.6 Reward function ablation
Given that the reward scale fluctuates with adjustments to these parameters, as indicated in Equations 8 and 7, we introduced the normalized score – recall-over-cost ratio as a metric for assessment of the results. For a comprehensive evaluation of accuracy, we utilized the BA score. The results are illustrated in Figure 7. Our findings revealed that the optimal parameter combination was $\alpha=0.5$ and $\beta=0.3c$. This combination yielded a BA score of $60\%$ and a recall-over-cost ratio of $67\%$, the highest scores among other combinations.
@@ 6 Discussion and conclusions
Given the complexity of KOA progression, it is worth discussing the various sensing policies we implemented as baselines. The annual sensing, while thorough, may not be cost-effective as it incurs the highest acquisition costs. The BAS policy seems to be an ideal easy-to-implement policy cost-wise, but it detects only $50\%$ of progressors. Implementing predictive models as a base for AS was expected to bring positive outcomes, however, the results did not yield a positive reward in any of the cases. In a general cost setting over $4$ years, to collect over $70\%$ of all progression events, the acquisition budget requires \$$\$1,400-\$2,000$ for each participant, whereas our policy costs only around \$$\$1,110$, which would help to save \$$\$300-\$700$ (approximately $21\%-45\%$). The interpretation of these numbers should be made from the budgetary perspective of a data collection organizer, which could be a hospital or a pharmaceutical company. The flexibility in various financial factors is an outstanding advantage of our method compared to any method we compared it to. The quality of decision-making of our policy depends on the ratio of acquisition cost over treatment cost ($\frac{\lambda}{c}$). We found that all settings where $\frac{\lambda}{c}\leq 0.3$ are the ones giving positive rewards. This threshold may assist the clinical trial or experiment organizer in budgetary planning, directly connecting decision-making and costs.
@@
We showed awareness of our policy regarding KOA characteristics such as disease severity, participant weight, and knee symptoms through more frequent follow-ups for patients at higher risk of OA. Interestingly, females are generally considered to be at higher risk of developing KOA than males [53, 14], albeit our policy selected more follow-ups for males. This observation can be partially explained by data bias in our test set. $30\%$ of females were healthy in terms of BMI, while the ratio of males with healthy BMI was only $15\%$. A similar observation was done for KLG subgroups, where $32\%$ of females but only $24\%$ of males were diagnosed as healthy at the baseline visit. Therefore, our policy arguably prioritized BMI over sex in the decision making. Interestingly, our policy tried to equalize the follow-up chance for everyone by recommending at least one follow-up even in the high visit cost scenario. Equalizing decisions can be considered as a drawback from the viewpoint of the data collection organizer. Lastly, our reward function did not implement a group-centered adaptive mechanism to maximize the efficacy in clinical trials (e.g. for prioritizing the regular visits for younger or more severe patients), which we will leave for future work.
@@
Our study is novel in proposing RL-based AS for KOA management. While we acknowledge certain areas for improvement, they also represent exciting opportunities for future research and development. Firstly, our method requires data from a longitudinal cohort for training. The scarcity of such data limits our ability to test with different populations. Secondly, our policy adaptively changes in various cost settings, and it might not always yield positive outcomes, especially in case when data acquisition cost is high. The presence of a “trade-off” point can be viewed as an upper bound of the monetary efficiency of an AS policy for KOA. Optimizing this bound will not only improve our model but will naturally results in the new knowledge in the field of KOA. Thirdly, we recognize the potential of tracking also symptomatic biomarkers when assessing disease trajectory. As KOA often causes chronic pain, integrating imaging biomarkers with symptomatic ones might lead to more precise progression outcome and, thus, more accurate progression modeling. Fourthly, our workflow used a pre-trained DNN model to predict KLG and extract probability for state input. While the performance of the pre-trained model can impact our results, it also highlights the importance of continual model refinement and improvement. However, for a RL workflow, working with high-dimensional data, such as plain radiographs, as the input state poses computational challenges. As we aimed to deliver a simple end-to-end AS solution in this work, the task of seeking more efficient and effective models is our motivation in the future. Consequently, while we acknowledge the challenges related to data limitations, we view these not as obstacles but as motivations for innovation and improvement in our method.
@@
To summarize, we have proposed a novel way to collect data for osteoarthritis trials utilizing RL. Our method requires no human input and can run fully automatically. We demonstrated that the proposed method outperforms existing baselines, and we evaluated different design choices through a set of comprehensive ablation studies. The developed approach to acquiring the next-generation data in OA shows promise in delivering more cost-effective clinical trials and observational studies in OA. Furthermore, it may also find applications in development of novel OA screening and/or rehabilitation programs. We believe that our work makes a significant step in accelerating the market introduction of OA DMOADs. To advance further development of AS in OA, all our codes are made publicly available at https://github.com/Oulu-IMEDS/OACostSensitivityRL.
@@
    Random. Randomly chooses an action at every time point for every patient.

### v2

@@ 3.1 Knee osteoarthritis
In the KOA research area, there has been a growing interest in the application of machine learning (ML) methods for KOA progression prediction [74, 77, 23, 43, 32, 76, 51, 49, 54, 55, 13, 27, 55]. Two main lines of work related to the prediction of KOA progression from clinical and imaging data are the prediction of TKR [77, 43, 32] and prediction of radiographic KOA worsening [74, 23, 49].
@@
The main limitation of most previous studies on KOA progression prediction is that they consider predictions per knee, disregarding that OA often co-occurs in two knees. The second limitation is that they ignore the long disease progression timelines, taking for modeling only one baseline measurement point. OA can progress very slowly, and be drastically accelerated by various future unknown factors, e.g. injury [poulsen2019knee]. Therefore, we believe that it is crucial to update the model’s predictions when newer data from a patient becomes available. Finally, none of the previous studies considered the financial aspect of their results, nor downstream drug development or clinical applications. Our work aims to address all these limitations, enabling dynamic patient-level data collection, while balancing data collection utility and costs with RL.
@@ 3.2 Active sensing
RL has become one of the most common techniques to explore AS in recent years, particularly in medical applications for disease management [2, 83, 10, 62, 29]. DPSCREEN framework proposed by [2] aimed to deliver a personalized screening policy for breast cancer (BC) patients. The proposed policy takes into account both patient features and their clinical history, and it demonstrated significant reductions in the number of screenings performed while maintaining the same expected delays in disease detection. A similar problem of BC screening was also later tackled by [81].
[83] performed the AS using the Actor-Critic (ASAC) framework to tackle the issue of selecting the variables that should be observed. The ASAC framework includes two networks: a selector network, playing as an actor, and a predictor network, playing as a critic. The authors showed that ASAC outperformed state-of-the-art methods in two real-world medical datasets of Alzheimer’s Disease and electronic health records.
[10] presented a deep Q-learning approach to strategically schedule measurements, especially in ICU mortality prediction. In simulations and real-world datasets, the proposed policy outperforms heuristic-based scheduling with higher predictive gain and lower cost.
The prior work on AS with RL focused on improving general RL methods and placed a substantially lesser focus on the reward function, the disease of interest, as well as the challenge of handling high-dimensional data. In this work, we chose a simple and the most common RL technique – Q-learning and focused on the development of the reward and the environment for the target domain of KOA.
@@ 4.2 Hyper-parameters of the reward function
We empirically chose $\alpha=0.5$ as a parameter to scale down the late visit-associated negative reward and set $\beta=0.3c$ to assign positive rewards for correct dismissal. $\lambda=0.5$ was used in all our experiments, and we also conducted an ablation study to validate how our method performs in different settings. More details about these hyper-parameters will be presented in Section 5.2.
@@ 4.3 Agent
[…]
4.4 Evaluation
We compared our method to a wide range of reference methods belonging to three categories of sensing – random, routine, and active sensing. Random policies combined $4$ probabilities of taking follow-up action: $0.3$, $0.5$, $0.7$, and $0.9$. The set of routine policies comprised “no sensing” (NS), annual sensing (ANS), and bi-annual sensing (BAS). The active policies set included logistic regression (LR), cox regression (CR), RNN-based methods, and CLIMATv2 [51, 49]. The reference methods are described in detail in Section S1.1.
@@ 5.2 Impact of financial factors on decision-making
We demonstrate the results of the second experiment in Figure 4b. The performance of CLIMATv2 and the NS policy exhibited a decreasing trend as the JSW cost, $c$, increased while keeping unchanged $\lambda=0.5$. In contrast, we observed that the ANS policy generally increased in proportion to $c$. Our policy achieved a constant reward of $-0.25$ for $c\leq 1.0$, and demonstrated a linear increase for $c>1.0$. Overall, our proposed policy gained the highest reward among the methods at any value of $c\in[0.5,2]$. For values of $c$ greater than $1.5$, our method stands out as the sole approach yielding positive rewards per person.
@@ 5.3 Patient-level versus knee-level predictions
Table 4 demonstrates the overall performance of CLIMATv2 and our proposed method with knee-level and patient-level approaches. Notably, the patient-level approach outperformed the knee-level approach for both CLIMATv2 and the RL-based policy, with average improvements in BA of $1.44\%$ and $3.73\%$, respectively, and in RC of $4.58\%$ and $20.25\%$, respectively. Both the methods gained a $0.02$ increase in RPP when switching from the knee-level to the patient-level approach.
To provide additional insights, we further present the corresponding fine-grained results in Figure 5. Most of the time, patient-level models reduced the standard error in BA and RC of both CLIMATv2 and our RL-based agent. Particularly, at the 1-year time point, our patient-level policy exceeded the patient-level CLIMATv2 with the highest difference by approximately 15% in BA and 60% in RC. Furthermore, the greatest gap between patient-level and knee-level of our method occurred 3 years after baseline by approximately 20% in BA and 40% in RC. In general, our patient-level policy showed substantial improvement compared to other methods most of the time.
@@ 5.4 Importance of imaging data
We also evaluated the impact of imaging data on the performance of the proposed method. To set up a benchmark, we conducted experiments without the contributions of radiographs, thereby excluding the KLG probabilities from the agent’s input. We then incorporated one of the three distinct types of KLGs into our experiments. The first type, referred to as the “KLG ground truth” corresponds to the manual grades produced by the radiologists. The second and third types named the “image features” and the “KLG probabilities”, respectively, were extracted from a DNN that had been trained to predict the KLGs from radiographs automatically [tiulpin2018automatic, 50].
@@ 5.6 Reward function ablation
Given that the reward scale fluctuates with adjustments to these parameters, as indicated in Equations 13 and 12, we introduced the normalized score – recall-over-cost ratio as a metric for assessment of the results. For a comprehensive evaluation of accuracy, we utilized the BA score. The results are illustrated in Figure 7. Our findings revealed that the optimal parameter combination was $\alpha=0.5$ and $\beta=0.3c$. This combination yielded a BA score of $60\%$ and a recall-over-cost ratio of $67\%$, the highest scores among other combinations.
@@ 6 Discussion and conclusions
Given the complexity of KOA progression, it is worth discussing the various sensing policies we implemented as baselines. The annual sensing, while thorough, may not be cost-effective as it incurs the highest acquisition costs. The BAS policy seems to be an ideal easy-to-implement policy cost-wise, but it detects only $50\%$ of progressors. Implementing predictive models as a base for AS was expected to bring positive outcomes, however, the results did not yield a positive reward in any of the cases. In a general cost setting over $4$ years, to collect over $70\%$ of all progression events, the acquisition budget requires \$$\$1,400-\$2,000$ for each participant, whereas our policy costs only around \$$\$1,110$, which would help to save approximately $21\%-45\%$ of costs per patient in a trial. The flexibility in various financial factors is an outstanding advantage of our method compared to any method we compared it to. The quality of decision-making of our policy depends on the ratio of acquisition cost over treatment cost ($\lambda/c$). We found that all settings where $\lambda/c\leq 0.3$ are the ones giving positive rewards. This threshold may assist the clinical trial or experiment organizer in budgetary planning, directly connecting decision-making and costs.
@@
We showed awareness of our policy regarding KOA characteristics such as disease severity, participant weight, and knee symptoms through more frequent follow-ups for patients at higher risk of OA. Interestingly, females are generally considered to be at higher risk of developing KOA than males [53, 14], albeit our policy selected more follow-ups for males. This observation can be partially explained by data bias in our test set. $30\%$ of females were healthy in terms of BMI, while the ratio of males with healthy BMI was only $15\%$. A similar observation was done for KLG subgroups, where $32\%$ of females but only $24\%$ of males were diagnosed as healthy at the baseline visit. Therefore, as a risk factor, our policy prioritized BMI over sex in the decision-making. Interestingly, our policy tried to equalize the follow-up chance for everyone by recommending at least one follow-up even in the high visit cost scenario. Equalizing decisions can be considered as a drawback from the viewpoint of the data collection organizer.
@@
Our study’s core novelty is in proposing RL-based AS for KOA progression monitoring. While we acknowledge certain areas for improvement, they also represent exciting opportunities for future research and development. Firstly, our method requires data from a longitudinal cohort for training. The scarcity of such data limits our ability to test with different populations. On the other hand, we were able to utilize openly available OAI cohorts, and thus think that the developed policies can be fine-tuned on specific data. Secondly, our policy adaptively changes in various cost settings, and it might not always yield positive outcomes, especially in cases when data acquisition cost is high. The presence of a “trade-off” point can be viewed as an upper bound of the monetary efficiency of an AS policy for KOA. Optimizing this bound will not only improve our model but will naturally result in new knowledge in the field of KOA. Thirdly, we recognize the potential of tracking also symptomatic biomarkers when assessing disease trajectory. As KOA often causes chronic pain, integrating imaging biomarkers with symptomatic ones might lead to more precise progression outcomes and, thus, more accurate progression modeling. Fourthly, our workflow used a pre-trained DNN model to predict KLG and extract probability for state input. While the performance of the pre-trained model can impact our results, it also highlights the importance of continual model refinement and improvement. However, for an RL workflow, working with high-dimensional data, such as plain radiographs, as the input state poses computational challenges. While we acknowledge the mentioned limitations, we view these not as obstacles but as motivations for innovation and improvement in our method.
@@
To summarize, we have proposed a novel way to collect data for osteoarthritis trials utilizing RL. Our method requires no human input and can run fully automatically. We demonstrated that the proposed method outperforms existing baselines, and we evaluated different design choices through a set of comprehensive ablation studies. The developed approach to acquiring the next-generation data in OA shows promise in delivering more cost-effective clinical trials and observational studies in OA. Furthermore, our method may also find applications in the development of novel OA screening and/or rehabilitation programs, or translate to other domains. We believe that our work makes a significant step in accelerating the market introduction of OA DMOADs. To advance further development of AS in OA, all our codes are made publicly available at https://github.com/Oulu-IMEDS/OACostSensitivityRL.
@@
    Random. Randomly chooses an action with different probabilities at every time point for every patient.